given f(x)=4x-8 amd g(x)=2x-5 find the following
a. f(-1) b. g(-2) c. F(2) + g (-4)

Answer by 983manoj@gmail.com(4)

f(x)=4x-8 g(x)=2x-5

a.f(-1)=4(-1)-8
=-4-8
=-12
b.g(-2)=2(-2)-5
=-4-5
=-9
c.f(2)=4*2-8
=0
d.g(-4)=2(-4)-5
=-8-5
=-13